OpenStudy (anonymous):

The Volume V=(4/3)(pi)r^3 of a spherical balloon changes with the radius. At what rate does the volume change with respect to the radius when r=2ft? I know I have to take the derivative of V to get the answer, but why is the derivative of V 4(pi)r^2? I thought that the derivative of a constant was 0, so shouldn't it be 0(0)r^2, which is 0?

OpenStudy (anonymous):

Does that make sense at all?

OpenStudy (anonymous):

\[\frac{ dv }{ dr }=3*\frac{ 4 }{ 3 } \pi r^2=4 \pi r^2\] when r=2 \[\frac{ dv }{ dr }=4 \pi *2^2=16 \pi\] here neither v nor r is constant

OpenStudy (anonymous):

But 4/3 and pi are both constants, and the derivative of both of those are 0?

OpenStudy (anonymous):

Oh wait, is that only when they are being added and not multiplied?

OpenStudy (phi):

example: d/dx (3x)= 3 d/dx (x) = 3 the constant is "factored out"

OpenStudy (anonymous):

Oh, so the (4/3)pi is treated as the coefficient?

OpenStudy (anonymous):

So its the same as if it were (4pi)/3

OpenStudy (anonymous):

\[\frac{ d }{ dx }\left( a x^n \right)=a~n x ^{n-1},where ~a~ is ~a ~constant.\]

OpenStudy (anonymous):

And because you multiply the coefficient by the power, the 3 becomes 1. Ooooooh, that makes more sense now

OpenStudy (phi):

yes. If we went back to the definition of "derivative", (which is a pain) you can see that constant coefficients are factored out.

OpenStudy (anonymous):

Oh, ok! Thanks a lot guys!!!

OpenStudy (phi):

btw, Paul's Online Notes lists almost all the variations on "related rates" problems http://tutorial.math.lamar.edu/Classes/CalcI/RelatedRates.aspx